# NaturalReasoning: Reasoning in the Wild with 2.8M Challenging Questions

### v3

@@ 1 Introduction
Large language models (LLMs) have demonstrated increased reasoning capabilities (OpenAI et al., 2024; Guo et al., 2025). These models are designed to devote more time to deliberation before responding, enabling them to tackle intricate tasks and solve more complex problems in science, coding, and mathematics. Such reasoning models are trained via large-scale reinforcement learning on tasks where the reward can be derived using rule-based verification. Existing reasoning datasets are often limited to narrow domains where the solutions are short and easy to verify, while the majority of reasoning problems across broader domains are open-ended reasoning. To bridge this gap, we introduce NaturalReasoning, a comprehensive dataset curated from pretraining corpora, comprising 2.8 million reasoning questions spanning various topics, including Mathematics, Physics, Computer Science, Economics & Business, etc. NaturalReasoning is compared to a wide range of reasoning datasets, showcasing its advantageous properties, in particular its diversity and difficulty.
NaturalReasoning possesses several desirable attributes as a dataset, serving multiple research purposes. Firstly, the questions are backtranslated from the pretraining corpora, ensuring that it represents diverse reasoning problems in the real world, as opposed to synthetic datasets derived from benchmark datasets like MetaMathQA (Yu et al., ) and OpenMathInstruct-2 (Toshniwal et al., ). Secondly, it consists of both questions with easy-to-verify answers and those with open-ended solutions (e.g., theorem proving), providing a rich resource for developing learning algorithms to enhance LLMs’ reasoning abilities across broader domains.Thirdly, we show that NaturalReasoning poses more difficult reasoning problems than existing datasets. Its questions therefore provide an effective testbed for improving LLM reasoning—whether through knowledge distillation from a stronger teacher model or reinforcement learning with external and self-generated reward signals (Yuan et al., ). Lastly, the NaturalReasoning dataset complements existing reasoning datasets in terms of both quality and quantity.
@@ 2 Data Collection
Backtranslating questions based on pretraining corpora has been shown to be a scalable approach to create instruction-following datasets (Li et al., 2023; Yue et al., 2024). We take a similar approach to extract diverse and realistic reasoning questions grounded in pretraining data, i.e. we generate grounded synthetic data. An overview of the data creation approach is illustrated in Figure 1. A key differentiator of our approach is its emphasis on simplicity; we use LLMs throughout the entire cycle of synthetic data generation and curation, without any human annotation nor manual steps such as preprocessing to detect relevant documents and extracting questions with rule-based filtering.
@@ 2.1 Annotate Reasoning
Inspired by the meta-cognitive capabilities of state-of-the-art LLMs (Didolkar et al., 2024), we use an LLM to perform detailed annotation of pretraining documents to detect documents which contain sophisticated reasoning traces. We use the public pretraining corpora DCLM-baseline  (Li et al., 2024b) and FineMath  (Allal et al., 2025) as sources, which have demonstrated steeper scaling laws than alternative corpora. More specifically, given a document $d$ from the pretraining corpus, we prompt an LLM to rate the content in $d$ along multiple axes: Problem Completeness, Problem Complexity and Technical Depth, Technical Correctness and Accuracy, Thinking and Reasoning. The detailed prompt is provided in Appendix G. Empirically, we found that the model could analyze the document well and is able to follow the format requirement in the instruction.
@@
MetaMathQA Math 395K 41$\pm$24 Rephrased by GPT-3.5-Turbo from GSM8K+MATH GPT-3.5-Turbo
NuminaMath Math 860K 48$\pm$32 Grade-level and competition-level math datasets GPT-4o
OpenMath2 Math 607K 46$\pm$44 Synthesized by Llama3-405B-Inst from GSM8K+MATH Llama3-405B-Inst
WebInstruct Multi 13M 34$\pm$28 Recall and Extracted from Web Mixtral-22B×8, Qwen-72B
NaturalReasoning Multi 2.8M 55$\pm$21 Synthesized by Llama-3.3-70B-Inst grounded in Web Llama-3.3-70B-Inst
@@ 2.2 Synthesize Questions and Answers
For documents which are identified with a high degree of reasoning (e.g., scored full points on axes of Problem Complexity and Technical Depth, Thinking and Reasoning), we further prompt an LLM to compose a self-contained and challenging reasoning question $q$ based on the content in $d$. Different from existing work, which extracts questions appearing in the document (Yue et al., 2024), our approach allows us to synthesize more novel questions not directly contained in pretraining corpora. Then, we prompt the LLM to verify whether a correct reference answer $a$ to the synthesized question $q$ can be derived from $d$ and, if possible, include it as a reference answer. Finally, for every question we generate an additional response with a relatively strong open-source model (Llama-3-70B-Instruct), which we later use as a teacher signal for knowledge distillation (see Section 5).
@@ 2.3 Question Deduplication and Decontamination
  • Decontamination We filter out questions that are similar to popular reasoning benchmarks including MATH (Hendrycks et al., 2021b), GPQA, MMLU-Pro (Wang et al., 2024) and MMLU-Stem (Hendrycks et al., 2021a). The standard 13-gram decontamination method from EleutherAI’s llm-eval-harness (Gao et al., 2024) is used to identify and remove $0.026\%$ items from the dataset.
@@ 3 Data Analysis
  • MetaMathQA is created by bootstrapping mathematical questions from GSM8K and MATH, rewriting them from multiple perspectives to form a new dataset (Yu et al., 2024). The responses are generated using GPT-3.5-Turbo.
  • NuminaMath is a comprehensive collection of 860K pairs of math problems and solutions (Li et al., 2024c). The questions cover multiple sources including grade-level questions and competition problems. The solutions in NuminaMath dataset are generated or rewritten by GPT-4o.
  • OpenMathInstruct-2 is a collection of 14M synthesized math questions and solutions, based on GSM8K and MATH (Toshniwal et al., 2024). The solutions are generated by Llama3.1-405B-Instruct (Grattafiori et al., 2024) and curated through majority vote on the final answer.
  • WebInstruct recalls relevant documents from Common Crawl using a fastText model trained on a diverse seed dataset of quiz websites. It then extracts question-answer pairs contained in recalled web pages and uses LLMs (Qwen-72B (Bai et al., 2023), Mixtral-8×7B (Jiang et al., 2024)) to refine the answer (Yue et al., 2024).
@@ 3.3 Question Diversity
We use an off-the-shell sentence encoder [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to generate embeddings for questions in WebInstruct and NaturalReasoning. We then apply UMAP (McInnes et al., 2018) to project the high-dimensional embeddings into a 2D space, followed by K-means clustering (Wu, 2012) to identify distinct groups. We use Mixtral-8B to assign high-level labels to these clusters, which is prompted with a few examples from each cluster. As Figure 5 shows, NaturalReasoning contains a more diverse and dense representation of non-mathematical topics, including Physics, Chemistry, Computer Science, and Law, besides Math. In contrast, WebInstruct is primarily skewed toward mathematical content, highlighting the broader topic coverage of NaturalReasoning.
@@ 3.4 Reference Answer Analysis
We provide examples of questions with single-word, short, and long answers in Appendix B. In general, we found that questions with single word answers typically involve numerical, factual, or definitional queries, while questions with long answers demand more free-form in-depth analysis. For questions with a long answer, the extracted reference answer is typically a short summary content from the original documents or useful clues to answer the question. While reference answers may contain some noise, we demonstrate their potential usefulness in Appendix F.
@@ 4 Experimental Setup
We highlight the efficacy of NaturalReasoning in two settings: (1) Knowledge distillation, and (2) Unsupervised Self-Training. For (1), we evaluate whether NaturalReasoning enables steeper scaling than existing datasets when distilling reasoning capabilities to a student model via supervised finetuning (Section 5. We experiment with different model families such as Llama3.1-8B and Qwen2.5-7B. Specifically we show that questions from NaturalReasoning are very effective at distilling long chain-of-thoughts from reasoning models and compare it to manually curated questions such as LIMO (Ye et al., 2025) and S1K (Muennighoff et al., 2025) (Section 6). To demonstrate (2), we evaluate how well NaturalReasoning supports self-training either through a strong external reward model or self-rewarding mechanisms (Yuan et al., 2024) (Section 7).
@@
We evaluate our models on a diverse set of benchmarks that encompass both math and science reasoning: MATH, GPQA, GPQA-Diamond (Rein et al., 2024), and MMLU-Pro. To ensure a fair and consistent comparison, we adopt a zero-shot evaluation setting across all trained models. For inference we use vllm (Kwon et al., 2023) and employ greedy decoding to maintain determinism and eliminate variability introduced by stochastic generation. Unless mentioned otherwise, we report accuracy averaged over the last three saved model checkpoints during training.
@@ 5 Steeper Scaling with Challenging and Diverse questions
We fine-tuned the Llama3.1-8B-Base model and Qwen2.5-7B using fairseq2 training recipes (Balioglu, 2023), exploring the impact of varying dataset sizes. Specifically, we trained models on our dataset and the comparison datasets introduced in Table 1, and evaluated their average performance across three benchmarks: MATH, GPQA, and MMLU-Pro. For all datasets, we train 3 epochs for each data size, using batch size 128, learning rate $5e^{-6}$, with cosine learning rate schedule where final learning rate is 1% of peak learning rate.
@@ 5.1 Results
As shown in 4(a), models trained on NaturalReasoning require fewer training examples to achieve superior performance. With just 1.5 million training examples, the model trained on NaturalReasoning already outperforms Llama3.1-8B-Instruct, which was extensively tuned for instruction-following with more data (Grattafiori et al., 2024). In contrast, other datasets, including OpenMathInstruct-2 and WebInstruct, fail to surpass Llama3.1-8B-Instruct even when trained on 2.8 million data. Each NaturalReasoning sample therefore provides denser, more effective reasoning supervision, making it the most data-efficient choice for boosting model reasoning performance.
@@ 6 Eliciting Long Chain-of-Thought
In addition to the emergence of OpenAI-o1 and DeepSeek-R1, several studies suggest that simpler tasks require fewer steps while complex tasks benefit significantly from longer CoTs (Jin et al., 2024). When encouraging the model to think for longer, they are able to solve questions that they previously could not. Motivated by this, we investigate whether the questions in NaturalReasoning are complex enough to benefit from longer CoTs from a stronger reasoning model. We do so by distilling Deepseek-R1 responses to Llama3.3-70B-Instruct.
We randomly sampled 1K questions from NaturalReasoning and used SGLang (Zheng et al., 2023) to prompt DeepSeek-R1, generating one response per question. Resulting response lengths range from 745 to 14.6K tokens with an average length of 4430 tokens. We then supervised finetune Llama-3.3-70B-Instruct on this set and compare its performance against training on two strong, heavily curated datasets: s1K-1.1 (Muennighoff et al., 2025) and LIMO (Ye et al., 2025). Both datasets underwent multiple filtering stages to ensure that their questions are high-quality, diverse, and challenging; for consistency, all responses were generated with DeepSeek-R1. To keep the evaluation consistent with the setting used in Guo et al. (2025), we report pass@1 averaged across $n=24$ samples. Each sample is generated using temperature=$0.6$, top_p=$0.95$.
@@
Table 2: Pass@1 of Llama-3.3-70B-Instruct after distilling DeepSeek-R1 responses. We compare the performance of random selection from NaturalReasoning with curated datasets such as s1K-1.1(Muennighoff et al., 2025) and LIMO(Ye et al., 2025), as well as the scaling effect of NaturalReasoning.
@@ 7 Unsupervised Self-Training
Since open-ended reasoning questions are difficult to evaluate using exact match with reference answers, we explore whether our dataset can facilitate self-training through either strong external reward models or self-reward mechanisms (Yuan et al., )
@@
We verify unsupervised self-training under two different training method: Rejection-based sampling Fine-Tuning (RFT) and Direct Preference Optimization (DPO) (Rafailov et al., 2023), focusing on the effectiveness of different reward scoring strategies. Each approach relies on sampling 32 candidate responses per question, followed by selecting responses based on reward scores. RFT employs rejection sampling, selecting the highest-scoring response for SFT training, while DPO constructs training pairs using both the highest and lowest-scoring responses. For external reward models, we consider Qwen2.5-Math-RM-72B (Yang et al., 2024) and INF-ORM-Llama3.1-70B [^3] https://huggingface.co/infly/INF-ORM-Llama3.1-70B. In addition, we explore a self-rewarding framework where the model evaluates and assigns rewards to its own generated responses. Specifically, we consider the following self-rewarding strategies:
@@ 8 Related Work
Synthetic data has emerged as a promising solution for improving performance. Some approaches bootstrap new data from existing data (e.g., STaR (Zelikman et al., 2022) augments with new CoT rationales and MetaMath (Yu et al., ) rewrites the questions in MATH and GSM8K in several ways), but these techniques rely on the existence of a high-quality dataset. Other techniques such as that of OpenMathInstruct-2 (Toshniwal et al., ), Xwin-Math (Li et al., 2024a), and Self-Instruct (Wang et al., 2023) generate new data from only a few seed examples using an LLM but scaling to new domains remains a significant challenge. MMIQC (Liu et al., ) parses QA pairs from Mathematics Stack Exchange, using the highest-ranked answer, but few measures are taken to curate for quality and the resulting dataset is also specific to the math domain. Similar to our work, WebInstruct (Yue et al., 2024) harvests question-answer pairs from pre-training corpora and spans multiple domains, but is dependent on carefully crafted rule-based filters.
@@
Most prior works typically depend on human-annotated (gold) final answers (Zelikman et al., 2022; Chen et al., 2024; Pang et al., 2024) or the use of an external reward model (Singh et al., 2024; Dong et al., 2023). However, manually annotating or verifying final answers is particularly resource-intensive for complex, multi-step problems and training effective reward models for reasoning often requires human evaluation of LLM outputs (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023), making it similarly costly. Like works such as She et al. (2024); Yuan et al. (2024); Rosset et al. (2024); Tran et al. (2023), our work explores self-training in the absence of gold labels and does not limit itself to questions with short, easily verifiable targets.

### v4

@@ 1 Introduction
Large language models (LLMs) have demonstrated increased reasoning capabilities (openai2024openaio1card; guo2025deepseek). These models are designed to devote more time to deliberation before responding, enabling them to tackle intricate tasks and solve more complex problems in science, coding, and mathematics. Such reasoning models are trained via large-scale reinforcement learning on tasks where the reward can be derived using rule-based verification. Existing reasoning datasets are often limited to narrow domains where the solutions are short and easy to verify, while the majority of reasoning problems across broader domains are open-ended reasoning. To bridge this gap, we introduce NaturalReasoning, a comprehensive dataset curated from pretraining corpora, comprising 2.8 million reasoning questions spanning various topics, including Mathematics, Physics, Computer Science, Economics & Business, etc. NaturalReasoning is compared to a wide range of reasoning datasets, showcasing its advantageous properties, in particular its diversity and difficulty.
NaturalReasoning possesses several desirable attributes as a dataset, serving multiple research purposes. Firstly, the questions are backtranslated from the pretraining corpora, ensuring that it represents diverse reasoning problems in the real world, as opposed to synthetic datasets derived from benchmark datasets like MetaMathQA (yumetamath) and OpenMathInstruct-2 (toshniwal24openmathinstruct). Secondly, it consists of both questions with easy-to-verify answers and those with open-ended solutions (e.g., theorem proving), providing a rich resource for developing learning algorithms to enhance LLMs’ reasoning abilities across broader domains.Thirdly, we show that NaturalReasoning poses more difficult reasoning problems than existing datasets. Its questions therefore provide an effective testbed for improving LLM reasoning—whether through knowledge distillation from a stronger teacher model or reinforcement learning with external and self-generated reward signals (yuanself). Lastly, the NaturalReasoning dataset complements existing reasoning datasets in terms of both quality and quantity.
@@ 2 Data Collection
Backtranslating questions based on pretraining corpora has been shown to be a scalable approach to create instruction-following datasets (li2023self; yue2024mammoth2). We take a similar approach to extract diverse and realistic reasoning questions grounded in pretraining data, i.e. we generate grounded synthetic data. An overview of the data creation approach is illustrated in Figure 1. A key differentiator of our approach is its emphasis on simplicity; we use LLMs throughout the entire cycle of synthetic data generation and curation, without any human annotation nor manual steps such as preprocessing to detect relevant documents and extracting questions with rule-based filtering.
@@ 2.1 Annotate Reasoning
Inspired by the meta-cognitive capabilities of state-of-the-art LLMs (didolkar2024metacognitive), we use an LLM to perform detailed annotation of pretraining documents to detect documents which contain sophisticated reasoning traces. We use the public pretraining corpora DCLM-baseline (li2024datacomp) and FineMath (allal2025smollm2smolgoesbig) as sources, which have demonstrated steeper scaling laws than alternative corpora. More specifically, given a document $d$ from the pretraining corpus, we prompt an LLM to rate the content in $d$ along multiple axes: Problem Completeness, Problem Complexity and Technical Depth, Technical Correctness and Accuracy, Thinking and Reasoning. The detailed prompt is provided in Appendix I. Empirically, we found that the model could analyze the document well and is able to follow the format requirement in the instruction.
@@
[…]
@@ 2.2 Synthesize Questions and Answers
For documents which are identified with a high degree of reasoning (e.g., scored full points on axes of Problem Complexity and Technical Depth, Thinking and Reasoning), we further prompt an LLM to compose a self-contained and challenging reasoning question $q$ based on the content in $d$. Different from existing work, which extracts questions appearing in the document (yue2024mammoth2), our approach allows us to synthesize more novel questions not directly contained in pretraining corpora. Then, we prompt the LLM to verify whether a correct reference answer $a$ to the synthesized question $q$ can be derived from $d$ and, if possible, include it as a reference answer. Finally, for every question we generate an additional response with a relatively strong open-source model (Llama-3-70B-Instruct), which we later use as a teacher signal for knowledge distillation (see Section 5).
@@ 2.3 Question Deduplication and Decontamination
  • Decontamination We filter out questions that are similar to popular reasoning benchmarks including MATH (hendrycksmath2021), GPQA, MMLU-Pro (wang2024mmlu) and MMLU-Stem (hendryckstest2021). The standard 13-gram decontamination method from EleutherAI’s llm-eval-harness (eval-harness) is used to identify and remove $0.026\%$ items from the dataset.
@@ 3 Data Analysis
  • MetaMathQA is created by bootstrapping mathematical questions from GSM8K and MATH, rewriting them from multiple perspectives to form a new dataset (DBLP:conf/iclr/YuJSYLZKLWL24). The responses are generated using GPT-3.5-Turbo.
  • NuminaMath is a comprehensive collection of 860K pairs of math problems and solutions (numina_math_datasets). The questions cover multiple sources including grade-level questions and competition problems. The solutions in NuminaMath dataset are generated or rewritten by GPT-4o.
  • OpenMathInstruct-2 is a collection of 14M synthesized math questions and solutions, based on GSM8K and MATH (toshniwal2024openmath2). The solutions are generated by Llama3.1-405B-Instruct (grattafiori2024llama3herdmodels) and curated through majority vote on the final answer.
  • WebInstruct recalls relevant documents from Common Crawl using a fastText model trained on a diverse seed dataset of quiz websites. It then extracts question-answer pairs contained in recalled web pages and uses LLMs (Qwen-72B (bai2023qwentechnicalreport), Mixtral-8×7B (jiang2024mixtralexperts)) to refine the answer (yue2024mammoth2).
In addition, we compare models trained on NaturalReasoning with those trained on the OpenThoughts dataset (guha2025openthoughtsdatarecipesreasoning), a recent open-source collection designed for reasoning in math, code, and science. As shown in Appendix G, models trained on NaturalReasoning achieve better performance across general reasoning benchmarks, demonstrating its broader coverage and stronger generalization compared to OpenThoughts.
@@ 3.3 Question Diversity
We use an off-the-shell sentence encoder [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to generate embeddings for questions in WebInstruct and NaturalReasoning. We then apply UMAP (mcinnes2018umap) to project the high-dimensional embeddings into a 2D space, followed by K-means clustering (wu2012advances) to identify distinct groups. We use Mixtral-8B to assign high-level labels to these clusters, which is prompted with a few examples from each cluster. As Figure 5 shows, NaturalReasoning contains a more diverse and dense representation of non-mathematical topics, including Physics, Chemistry, Computer Science, and Law, besides Math. In contrast, WebInstruct is primarily skewed toward mathematical content, highlighting the broader topic coverage of NaturalReasoning.
@@ 3.4 Reference Answer Analysis
We provide examples of questions with single-word, short, and long answers in Appendix B. In general, we found that questions with single word answers typically involve numerical, factual, or definitional queries, while questions with long answers demand more free-form in-depth analysis. For questions with a long answer, the extracted reference answer is typically a short summary content from the original documents or useful clues to answer the question. While reference answers may contain some noise, we demonstrate their utility in Appendix F for both filtering training data in knowledge distillation and enabling reinforcement learning with verifiable rewards (RLVR) (lambert2025tulu3pushingfrontiers).
@@ 4 Experimental Setup
We highlight the efficacy of NaturalReasoning in two settings: (1) Knowledge distillation, and (2) Unsupervised Self-Training. For (1), we evaluate whether NaturalReasoning enables steeper scaling than existing datasets when distilling reasoning capabilities to a student model via supervised finetuning (Section 5. We experiment with different model families such as Llama3.1-8B and Qwen2.5-7B. Specifically we show that questions from NaturalReasoning are very effective at distilling long chain-of-thoughts from reasoning models and compare it to manually curated questions such as LIMO (ye2025limoreasoning) and S1K (muennighoff2025s1simpletesttimescaling) (Section 6). To demonstrate (2), we evaluate how well NaturalReasoning supports self-training either through a strong external reward model or self-rewarding mechanisms (yuan2024selfrewarding) (Section 7).
@@
We evaluate our models on a diverse set of benchmarks that encompass both math and science reasoning: MATH, GPQA, GPQA-Diamond (rein2024gpqa), and MMLU-Pro. In Appendix H, We also show NaturalReasoning’s utility for broader NLP tasks (e.g., writing). To ensure a fair and consistent comparison, we adopt a zero-shot evaluation setting across all trained models. For inference we use vllm (Kwon_2023) and employ greedy decoding to maintain determinism and eliminate variability introduced by stochastic generation. Unless mentioned otherwise, we report accuracy averaged over the last three saved model checkpoints during training.
@@ 5 Steeper Scaling with Challenging and Diverse questions
We fine-tuned the Llama3.1-8B-Base model and Qwen2.5-7B using fairseq2 training recipes (balioglu2023fairseq2), exploring the impact of varying dataset sizes. Specifically, we trained models on our dataset and the comparison datasets introduced in Table 1, and evaluated their average performance across three benchmarks: MATH, GPQA, and MMLU-Pro. For all datasets, we train 3 epochs for each data size, using batch size 128, learning rate $5e^{-6}$, with cosine learning rate schedule where final learning rate is 1% of peak learning rate.
@@ 5.1 Results
As shown in 4(a), models trained on NaturalReasoning require fewer training examples to achieve superior performance. With just 1.5 million training examples, the model trained on NaturalReasoning already outperforms Llama3.1-8B-Instruct, which was extensively tuned for instruction-following with more data (grattafiori2024llama3herdmodels). In contrast, other datasets, including OpenMathInstruct-2 and WebInstruct, fail to surpass Llama3.1-8B-Instruct even when trained on 2.8 million data. Each NaturalReasoning sample therefore provides denser, more effective reasoning supervision, making it the most data-efficient choice for boosting model reasoning performance.
@@ 6 Eliciting Long Chain-of-Thought
In addition to the emergence of OpenAI-o1 and DeepSeek-R1, several studies suggest that simpler tasks require fewer steps while complex tasks benefit significantly from longer CoTs (jin2024impact). When encouraging the model to think for longer, they are able to solve questions that they previously could not. Motivated by this, we investigate whether the questions in NaturalReasoning are complex enough to benefit from longer CoTs from a stronger reasoning model. We do so by distilling Deepseek-R1 responses to Llama3.3-70B-Instruct.
We randomly sampled 1K questions from NaturalReasoning and used SGLang (zheng2023sglang) to prompt DeepSeek-R1, generating one response per question. Resulting response lengths range from 745 to 14.6K tokens with an average length of 4430 tokens. We then supervised finetune Llama-3.3-70B-Instruct on this set and compare its performance against training on two strong, heavily curated datasets: s1K-1.1 (muennighoff2025s1simpletesttimescaling) and LIMO (ye2025limoreasoning). Both datasets underwent multiple filtering stages to ensure that their questions are high-quality, diverse, and challenging; for consistency, all responses were generated with DeepSeek-R1. To keep the evaluation consistent with the setting used in guo2025deepseek, we report pass@1 averaged across $n=24$ samples. Each sample is generated using temperature=$0.6$, top_p=$0.95$.
@@
Table 2: Pass@1 of Llama-3.3-70B-Instruct after distilling DeepSeek-R1 responses. We compare the performance of random selection from NaturalReasoning with curated datasets such as s1K-1.1(muennighoff2025s1simpletesttimescaling) and LIMO(ye2025limoreasoning), as well as the scaling effect of NaturalReasoning.
@@ 7 Unsupervised Self-Training
Since open-ended reasoning questions are difficult to evaluate using exact match with reference answers, we explore whether our dataset can facilitate self-training through either strong external reward models or self-reward mechanisms (yuanself)
@@
We verify unsupervised self-training under two different training method: Rejection-based sampling Fine-Tuning (RFT) and Direct Preference Optimization (DPO) (rafailov2023direct), focusing on the effectiveness of different reward scoring strategies. Each approach relies on sampling 32 candidate responses per question, followed by selecting responses based on reward scores. RFT employs rejection sampling, selecting the highest-scoring response for SFT training, while DPO constructs training pairs using both the highest and lowest-scoring responses. For external reward models, we consider Qwen2.5-Math-RM-72B (yang2024qwen25mathtechnicalreportmathematical) and INF-ORM-Llama3.1-70B [^3] https://huggingface.co/infly/INF-ORM-Llama3.1-70B. In addition, we explore a self-rewarding framework where the model evaluates and assigns rewards to its own generated responses. Specifically, we consider the following self-rewarding strategies:
@@ 8 Related Work
Synthetic data has emerged as a promising solution for improving performance. Some approaches bootstrap new data from existing data (e.g., STaR (zelikman2022star) augments with new CoT rationales and MetaMath (yumetamath) rewrites the questions in MATH and GSM8K in several ways), but these techniques rely on the existence of a high-quality dataset. Other techniques such as that of OpenMathInstruct-2 (toshniwal24openmathinstruct), Xwin-Math (li2024common), and Self-Instruct (wang2023self) generate new data from only a few seed examples using an LLM but scaling to new domains remains a significant challenge. MMIQC (liu2024augmenting) parses QA pairs from Mathematics Stack Exchange, using the highest-ranked answer, but few measures are taken to curate for quality and the resulting dataset is also specific to the math domain. Similar to our work, WebInstruct (yue2024mammoth2) harvests question-answer pairs from pre-training corpora and spans multiple domains, but is dependent on carefully crafted rule-based filters.
@@
Most prior works typically depend on human-annotated (gold) final answers (zelikman2022star; chen2024self; pang2024iterative) or the use of an external reward model (singh2023beyond; dong2023raft). However, manually annotating or verifying final answers is particularly resource-intensive for complex, multi-step problems and training effective reward models for reasoning often requires human evaluation of LLM outputs (cobbe2021training; uesato2022solving; lightman2023let), making it similarly costly. Like works such as she2024mapo; yuan2024selfrewarding; rosset2024direct; viethoangtranduong, our work explores self-training in the absence of gold labels and does not limit itself to questions with short, easily verifiable targets.